How do i use linear combinations to solve the
system of this linear equation
2a+6z=4
3a-7z=6

Answer by stanbon(75887)

2a+6z=4
3a-7z=6
-----
Multiply thru the 1st equation by 2.
Multiply thru the 2nd equation by 3.
---
6a + 18z = 12
6a - 14z = 12
---
Subtract and solve for "z":
32z = 0
z = 0
----
Solve for "a":
2a + 6z = 4
2a + 6*0 = 4
2a = 4
a = 2
------
Cheers,
Stan H.
==========